A parallel plate capacitor with plate separation o
Answer #1

(a) \varepsilonr =1, V =10V , d =1 mm

E =V/d =10/0.001 =104 V/m

Electric flux density D =\varepsilonr E = 1*104

D = 104 C/m2

(b) \varepsilonr =5, V =10V , d =1 mm

E =V/d =10/0.001 =104 V/m

Electric flux density D =\varepsilonr E = 5*104

D = 5*104 C/m2

Electric filed does not change. In E= V/d there isn no presence of dielectric

Electric flux density is changed. It depens on dielectric material.

(c) \varepsilonr =1, E = 50 kV/m

Electric flux density D =\varepsilonr E = 5*50*103

D = 2.5*105 C/m2

Electric filed does not change. In E= V/d there isn no presence of dielectric

Electric flux density is changed. It depens on dielectric material.